# Pragmatic Inference for Moral Reasoning Acquisition: Generalization via Distributional Semantics

### v1

@@ Abstract
Moral reasoning has emerged as a promising research direction for Large Language Models (LLMs), yet achieving generalization remains a central challenge. From a linguistic standpoint, this difficulty arises because LLMs are adept at capturing distributional semantics, which fundamentally differs from the morals which operate at the pragmatic level. This paper investigates how LLMs can achieve generalized moral reasoning despite their reliance on distributional semantics. We propose pragmatic inference methods grounded in moral foundations theory, which leverage contextual information at each step to bridge the pragmatic gap and guide LLMs in connecting moral foundations with moral reasoning objectives. Experimental results demonstrate that our approach significantly enhances LLMs’ generalization in moral reasoning, providing a foundation for future research grounded in moral foundations theory. Warning: this paper contains offensive language.
Pragmatic Inference for Moral Reasoning Acquisition: Generalization via Distributional Semantics
Guangliang Liu1    Xi Chen2    Bocheng Chen3    Xitong Zhang1  Kristen Johnson1 1Michigan State University  2Nanyang Technological University  3University of Mississippi {liuguan5,zhangxit,kristenj}@msu.edu     zoexi.chen@ntu.edu.sg    bchen5@olemiss.edu
@@ 1 Introduction
With the advancement of Large Language Models (LLMs), how to regulate their performance in accordance with human moral values has attracted scholarly interest. However, it still remains as a challenging task due to the gap between distributional semantics used by LLMs and the pragmatics-based nature of moral reasoning  Liu et al. (2025a); metaphorically speaking, the gap between what is said and what is morally implied. Distributional semantics follow Firth’s (1957) distributional hypothesis which argues that two linguistic units have similar meaning if they occur in similar contexts. In contrast, moral reasoning goes beyond this structuralism view of language and requires the capability of deriving conclusions from implicature of language use, moral norms and behavioral consequences, and observations of discursive rules of conduct Spencer-Oatey and Xing (2019). These unsaid variables need to be linguistically manifested so that the gap between distributional semantics and pragmatics in moral reasoning can be closed, which increases the methodological generalizability that previous studies have thus far struggled to achieve  Rao et al. (2023); Preniqi et al. (2024); Jiang et al. (2025)
@@
they get very scared? Reply: I don’t think I’ve ever
heard of that happening, but I guess it’s possible.
@@
Table 1: Moral Reasoning Example From Ziems et al. (2022). The Moral Judgment for the given reply is Agree, and the Rule of Thumb (RoT) represents a judgment regarding an action mentioned in the prompt-reply, grounded in the moral foundation of fairness.
Table 1 demonstrates an example of the several elements that are involved in moral reasoning: rule of thumb (RoT), moral foundation, and moral judgment Ziems et al. (2022). Thus far, learning paradigms, which aim to equip LLMs with moral reasoning, generally train models to directly map a situation (Prompt and Reply) to moral judgment, with or without a reference to moral foundations and/or RoT  Hendrycks et al. (2020); Forbes et al. (2020); Jiang et al. (2021); Ziems et al. (2022); Jiang et al. (2025). Direct mapping often requires a huge amount of data to enable LLMs to capture the statistical correlations between situations and moral reasoning objectives.
On the other hand, we have witnessed the effectiveness of pragmatic inference frameworks used in other morality-relevant tasks, such as the identification of social biases and toxicity  Sap et al. (2020); Zhou et al. (2023); Chen and Wang (2025). For example,  Sap et al. (2020) developed a structured inference path that enables the identification of stereotypical implications encoded in explicit toxic language. Chen and Wang (2025) developed an inference chain that provides LLMs with the freedom to interpret online neologisms and rephrase the literal meanings of an input. Their method improved the performance of LLMs in recognizing highly implicit toxic language by 12% to 20%. These approaches are consonant with our goal of bridging the gap between distributional semantics and pragmatics, for instance, through interpreting literal meanings Chen and Wang (2025).
However, moral reasoning requires different inferential methods from the aforementioned toxicity or bias identification. Unlike toxic language (e.g., lowering employment standards for females implying females are less capable), the moral foundations [^1] Please refer to Section 2 for more discussion on Moral Foundations Theory for moral reasoning. that underlie a moral situation are often not encoded by certain linguistic cues at all  Clark and Schaefer (1989); Graesser et al. (1994); Sap et al. (2022). Inferring them necessitates context-specific interpretations of abstract moral rules (e.g., the rule “do not hurt others” may manifest as “do not push another child to the floor“ in children’s playground and “do not emotionally abuse your partner“ for adult relationships). Therefore, abstract moral foundations, rules of conduct, and context all need to be sufficiently textualized so as to be incorporated in machines’ reasoning process.
@@
Experimental results show that our proposed pragmatic inference methods significantly improves generalization in moral reasoning tasks. Section 3 presents the benchmark annotation and experimental settings. Section 4 introduces our proposed pragmatic inference methods. Section 5 presents further analysis to characterize the achieved generalization with our proposed pragmatic inference methods. Finally, Section 6 concludes with a discussion of our findings and outlines potential directions for future work.
@@ 2 Related Works
Distributional Semantics is one of the theoretical foundations for embeddings. It interprets the meaning of a linguistic unit by additional signs (e.g., words, emoji) surrounding it Harris (1954); Firth (1957); McDonald and Ramscar (2001); Lenci and Sahlgren (2023); Boleda (2020). However, due to its nature as a structuralism approach to understanding language, distributional semantics suffers limitations. It treats the “context” purely as the texts provided, neglecting the many “unsaid” and latent parts of real interactions, such as the social constraints that a speaker encounters, the effects that a speech creates on the hearers, and the delivery of intention Lenci et al. (2008); Purver and Sadrzadeh (2015). These are studied in pragmatics, where LLMs have shown several deficiencies, e.g., understanding implicature Cong (2024), comprehending metaphors and humours Ruis et al. (2023); Barattieri di SanPietro et al. (2023), and moral alignment Khamassi et al. (2024); Liu et al. (2024).
@@
Pragmatic Inference and Moral Reasoning. In this work, we leverage pragmatic inference to facilitate LLMs to achieve stronger performance in moral reasoning. Pragmatic inference refers to the process of deriving conclusions of meanings based on context, intention, and language use Elder (2024). The inferential process is often not verbally realized and, instead, involves evaluations made at the metapragmatic level Kádár and Haugh (2013). That is, humans construct meta links that connect an instance of language use to certain social qualities (e.g., linking the phrase “thank you” to the concept of politeness). They learn these metapragmatic links over their physical and social experiences, which are entirely absent in LLMs. These limitations suggest that: (i) the inferential process, that is not verbally realized and often does not need to for humans, needs to be textualized so that LLMs can learn; and (ii) the context, where humans gain their experiences, need to sufficiently textualized, including both societal (e.g., social norms) and situational contexts (e.g., situated interpretations of language use). Without effective methods to textualize them, it is not a surprise that moral reasoning has long been a challenge in the broader field of artificial intelligence Allen et al. (2005); Van Rysewyk and Pontier (2014); Wallach et al. (2020).
@@
Moral Foundations Theory (MFT) Haidt and Joseph (2004); Graham et al. (2013) offers a framework for understanding human moral reasoning. It hypothesizes that moral judgments are guided by several domain-specific foundations, i.e., care, fairness, loyalty, authority, sanctity, and liberty. These foundations capture the diversity of moral concerns and have been widely adopted in psychology and computational social science Johnson and Goldwasser (2018). They provide a principled way to annotate, analyze, and model moral dimensions in language, making them particularly relevant for studying how LLMs represent and reason about morality Forbes et al. (2020); Ziems et al. (2022).
@@ 3 Preliminary
Benchmark. In this paper, we adopt the Moral Integration Corpus (MIC) benchmark and its annotations from Ziems et al. (2022). We chose the MIC benchmark because it closely reflects how humans make moral judgments in daily communication and provides richer contextual information. One advantage of the MIC dataset is that it provides levels of agreement between the annotators, which allows us to focus on the less controversial part of their data. Table 1 presents an example from MIC. The example consists of a Prompt (“Do people actually shit themselves when they get very scared??”) and a Reply (“I don’t think I’ve ever heard of that happening, but I guess it’s possible..”), with three evaluations annotated: agreement with the reply as moral judgment, RoT, and MF. The MIC benchmark utilizes the six MFs defined by Graham et al. (2013): care, fairness, liberty, loyalty, authority, and sanctity. We take two of the evaluations as our objectives: moral judgment (agree, neutral, or disagree with the reply) and MFs. RoT, instead, is used as part of the learning paradigm. RoT is not selected as a learning objective because: (i) it does not have a fixed answer and, more crucially, multiple RoTs may exist in one situation [^2] We refer readers to Section 3.1 of Ziems et al. (2022) for further details., and (ii) it has a rather dynamic relationship with MFs.
@@
Experimental Settings. Following the steps of our pragmatic inference frameworks (detailed in Section 4), we design several questions and prompt the deepseek-chat model [^3] https://api.deepseek.com to answer them. The step-wise answers are then used as the training data for fine-tuning the LLMs towards moral reasoning objectives. We chose Llama3-3.2 1B and 3B Grattafiori et al. (2024) as our backbone models, and leverage the supervised fine-tuning objective to train those models by following previous studies Ziems et al. (2022); Liu et al. (2025a). We intentionally select smaller models, as they are generally considered less capable than their larger counterparts. We aim to demonstrate the effectiveness of our pragmatic inference frameworks through the strong performance achieved by these smaller models.
@@
To reiterate, we have three pragmatic inference frameworks, corresponding to three tasks. For the moral judgment task (Agree, Neutral, Disagree with the reply), we use prediction accuracy as the evaluation metric. For the moral foundations classification (MFC) task, we compute prediction accuracy according to the number of MFs. Specifically, accuracies are reported separately for cases where a single RoT corresponds to one MF and for cases referring to two or more MFs. In other words, we have taken into consideration the one-to-many relationship between RoT/situation and MFs. Denoting the LLM $f$ is parameterized by $\theta$, the definition of MFs is represented with $d_{m}$ ((see Appendix A.1). Let $(x_{S},x_{R},y_{M},y_{J})$ denote a sample from the MIC benchmark, where $x_{S},x_{R}$ represents the moral situation and RoT, respectively. $y_{M}$ is its associated MF label. Each $y_{M}$ may contain between one and six MFs. $y_{J}$ is the moral judgment. We use $|y_{M}|=i$ $(1\leq i\leq 6,x\in\mathbb{Z})$ to indicate that $x$ is associated with $i$ MFs. We compute the accuracy for each value of $i$.
Input: There is a rule of thumb (RoT): {RoT}
There are the six MFs and their DEFINITIONS are:{DEFINITIONS}.
Inference: (1) The rule of thumb includes a Judgment to an Action, identify the Judgment and the
associated Action. (2) What would be the consequence if the Action is carried out? (3) Why is
the consequence of the Action relevant to the MFs {MFs} by referring to the provided
DEFINITIONS of those MFs.
@@
Table 2: Prompt Questions For the Task of MFs Classification. All content within { } is provided by the benchmark. Answer examples for the prompt questions are available in Appendix A.3.
@@
For baseline approaches, we follow prior work Forbes et al. (2020); Ziems et al. (2022) and consider two settings: (1) The base setting. For the MFC task, we compute $y_{J}=f_{\theta}(x_{S})$; and for moral judgment, we compute $y_{M}=f_{\theta}(x_{R})$. (2) The base+ setting. For the moral judgment task, we compute $y_{J}=f_{\theta}(x_{S},y_{M})$; and for MFC task, we compute $y_{M}=f_{\theta}(x_{R},d_{m})$. We additionally include the definitions of the six MFs ($d_{m}$) in the prompts. When the objective is moral judgment, we also explicitly provide the ground-truth MFs in the prompts. In other words, we intentionally provide the models with all available information in order to build strong baselines, which are then used to compare with the results of our framework. The setting of our proposed method is in line with the base+ setting but we incorporate an additional inference process in the input, e.g., $y_{J}=f_{\theta}(x_{S},y_{M},\{\text{inference}\})$.
@@
To examine how moral reasoning performance scales with the size of the training corpus, we consider three data settings. The benchmark reports agreement levels, and 26% of the samples receive full agreement between all annotators. We use subsets of 5K, 10K, and all the 23500 samples of the full-agreement data. All performance results are reported based on evaluation using the development set. More details about hyperparameters and experimental setup are in Appendix A.2.
@@ 4 Pragmatic Inference
In this section, we introduce our pragmatic inference methods for moral reasoning. Please note, pragmatic inferences often have more than one pathway to achieve a conclusion. Our approach represents one such pathway. Specifically, we build three inferential methods for moral foundation classification (MFC), the moral judgment, and MFC-judgment joint tasks, respectively.
@@ 4.1 Motivation
From the linguistic standpoint, our design of pragmatic inference frameworks is guided by three principles: (1) MFs are common ground which underlie all interactions. This has been supported by the theory and experiments of MFs Graham et al. (2013); Abdulhai et al. (2024). (2) MFs are abstract values extracted from discursive situations and hence need to be reified as context-specific interpretations Graham et al. (2011a). It is these context-specific interpretations of MFs which provide effective semantic cues for the models to reason morally. (3) Moral judgment is the judgment of good or bad by its nature, more specifically, the good/bad evaluation placed on the outcome of an action Nozick (1968); Schaich Borg et al. (2006); Emelin et al. (2021). Therefore, it is important to build the connections between contexts, discursive interpretations of MFs, and the outcome (i.e., behavioral consequence).
From a statistical standpoint, Liu et al. (2025a) demonstrate that the pragmatic gap between moral situations and moral reasoning objectives introduces greater degradation in language modeling performance than purely semantics-driven tasks, as LLMs tend to develop spurious associations that undermine distributional semantics. To address this, Moral Foundations Theory provides a principled signal: all moral situations can be interpreted through its lens. The key, then, is to uncover the linguistic cues that map moral situations to a unified hidden state space that is characterized by MFs, thereby alleviating the degradation of the language modeling performance.

### v2

@@ Abstract
Moral reasoning has emerged as a promising research direction for Large Language Models (LLMs), yet achieving generalization remains a central challenge. From a linguistic standpoint, this difficulty arises because LLMs are adept at capturing distributional semantics, which fundamentally differs from the morals which operate at the pragmatic level. This paper investigates how LLMs can achieve generalized moral reasoning despite their reliance on distributional semantics. We propose pragmatic inference methods grounded in moral foundations theory, which leverage contextual information at each step to bridge the pragmatic gap and guide LLMs in connecting moral foundations with moral reasoning objectives. Experimental results demonstrate that our approach significantly enhances LLMs’ generalization in moral reasoning, providing a foundation for future moral reasoning research based on pragmatic inference. Warning: this paper contains offensive language.
@@ 1 Introduction
With the advancement of Large Language Models (LLMs), how to enhance their capability of moral reasoning in order to regulate their performance in accordance with human moral values has attracted scholarly interest. However, it still remains as a challenging task due to the gap between distributional semantics used by LLMs and the pragmatics-based nature of moral reasoning  liu2025diagnosing; metaphorically speaking, the gap between what is said and what is morally implied. Distributional semantics follow Firth’s (firth1957synopsis) distributional hypothesis which argues that linguistic units have similar meaning if they occur in similar contexts. In contrast, moral reasoning goes beyond this structuralism view of language and requires the capability of deriving conclusions from implicature of language use, moral norms and behavioral consequences, and observations of discursive rules of conduct spencer-oatey_interdisciplinary_2019. These unsaid variables need to be linguistically manifested so that the gap between distributional semantics and pragmatics in moral reasoning can be closed, which increases the methodological generalizability that previous studies have thus far struggled to achieve rao2023ethical; preniqi2024moralbert; jiang2025investigating.
@@
they get very scared?
Reply: I don’t think I’ve ever heard of that
happening, but I guess it’s possible.
Rule of Thumb: It is wrong to shit on ourselves.
@@
Table 1: Moral Reasoning Example From ziems2022moral. The Moral Judgment for the given reply is Agree, and the Rule of Thumb (RoT) represents a judgment regarding an action mentioned in the prompt-reply, grounded in the moral foundation of fairness.
Table 1 demonstrates an example of the several elements that are involved in moral reasoning: rule of thumb (RoT), moral foundation, and moral judgment ziems2022moral. Thus far, learning paradigms, which aim to equip LLMs with moral reasoning, generally train models to directly map a situation (Prompt and Reply) to moral judgment, with or without a reference to moral foundations and/or RoT  hendrycks2020aligning; forbes2020social; jiang2021can; ziems2022moral; jiang2025investigating. Direct mapping often requires a huge amount of data to enable LLMs to capture the statistical correlations between situations and moral reasoning objectives.
On the other hand, we have witnessed the effectiveness of pragmatic inference used in other morality-relevant tasks, such as the identification of social biases and toxicity  sap2020social; zhou-etal-2023-cobra; chen2025pragmatic. For example,  sap2020social developed a structured inference path that enables the identification of stereotypical implications encoded in explicit toxic language. chen2025pragmatic developed an inference chain that provides LLMs with the freedom to interpret online neologisms and rephrase the literal meanings of an input. Their method improved the performance of LLMs in recognizing highly implicit toxic language by 12% to 20%. These approaches are consonant with our goal of bridging the gap between distributional semantics and pragmatics, for instance, through interpreting literal meanings chen2025pragmatic.
However, moral reasoning requires different inferential methods from the aforementioned toxicity or bias identification. Unlike toxic language (e.g., “lowering employment standards for females” implying “females are less capable”), the moral foundations***Please refer to Section 2 for more discussion on Moral Foundations Theory for moral reasoning. that underlie a moral situation are often not encoded by certain linguistic cues at all  clark1989contributing; graesser1994constructing; sap-etal-2022-neural. Inferring them necessitates context-specific interpretations of abstract moral rules (e.g., the rule “do not hurt others” may manifest as “do not push another child to the floor” on a children’s playground and “do not emotionally abuse your partner” for adult relationships). Therefore, abstract moral foundations, rules of conduct, and context all need to be sufficiently textualized so as to be incorporated in machines’ reasoning process.
@@
Experimental results show that our proposed pragmatic inference methods significantly improves generalization in moral reasoning tasks. Section 3 presents the benchmark annotation and experimental settings. Section 4 introduces our proposed pragmatic inference methods. Section 5 presents further analysis to characterize the achieved generalization with our proposed pragmatic inference methods. Finally, Section 7 concludes our findings and outlines potential directions for future work.
@@ 2 Related Works
Distributional Semantics is one of the theoretical foundations for embeddings. It interprets the meaning of a linguistic unit by additional signs (e.g., words, emoji) surrounding it harris1954distributional; firth1957synopsis; mcdonald2001testing; lenci2023distributional; boleda2020distributional. However, due to its nature as a structuralism approach to understanding language, distributional semantics suffers limitations. It treats the “context” purely as the texts provided, neglecting the many “unsaid” and latent parts of real interactions, such as the social constraints that a speaker encounters, the effects that a speech creates on the listeners, and the delivery of intention lenci2008distributional; purver2015distributional. These are studied in pragmatics, where LLMs have shown several deficiencies, e.g., understanding implicature cong_manner_2024, comprehending metaphors and humours ruis_goldilocks_2023; barattieri_di_sanpietro_pragmatic_2023, and moral alignment khamassi2024strong; liu2024towards.
@@
Pragmatic Inference and Moral Reasoning. In this work, we leverage pragmatic inference to facilitate LLMs to achieve stronger performance in moral reasoning. Pragmatic inference refers to the process of deriving conclusions of meanings based on context, intention, and language use elder_pragmatic_2024. The inferential process is often not verbally realized and, instead, involves evaluations made at the metapragmatic level kadar_understanding_2013. That is, humans construct meta links that connect an instance of language use to certain social qualities (e.g., linking the phrase “thank you” to the concept of politeness). They learn these metapragmatic links over their physical and social experiences, which are entirely absent in LLMs. These limitations suggest that: (i) the inferential process, that is not verbally realized and often does not need to be for humans, needs to be textualized so that LLMs can learn; and (ii) the context, where humans gain their experiences, needs to be sufficiently textualized, including both societal (e.g., social norms) and situational contexts (e.g., situated interpretations of language use). Without effective methods to textualize them, it is not a surprise that moral reasoning has long been a challenge in the broader field of artificial intelligence allen2005artificial; van2014hybrid; wallach2020machine.
@@
Moral Foundations Theory (MFT). haidt2004intuitive and graham2013moral describe a framework for understanding human morals and moral reasoning. MFT hypothesizes that moral judgments are guided by domain-specific foundations, i.e., care, fairness, loyalty, liberty, authority, and sanctity. These foundations capture the diversity of moral concerns and have been widely adopted in computational social science johnson2018classification; wendell2021classifying; reinig2024survey. They provide a principled way to annotate, analyze, and model moral dimensions in language, making them particularly relevant for studying how LLMs represent and reason about morality forbes2020social; ziems2022moral.
@@ 3 Preliminary
Benchmark. In this paper, we adopt the Moral Integration Corpus (MIC) benchmark and its annotations from ziems2022moral. We chose the MIC benchmark because it closely reflects how humans make moral judgments in daily communication and provides richer contextual information. One advantage of the MIC dataset is that it provides levels of agreement between the annotators, which allows us to focus on the less controversial part of their data. Table 1 presents an example from MIC. The example consists of a Prompt (“Do people actually shit themselves when they get very scared??”) and a Reply (“I don’t think I’ve ever heard of that happening, but I guess it’s possible..”), with three evaluations annotated: agreement with the reply as moral judgment, RoT, and MF. The MIC benchmark utilizes the six MFs defined by graham2013moral: care, fairness, liberty, loyalty, authority, and sanctity. We take two of the evaluations as our objectives: moral judgment (agree, neutral, or disagree with the reply) and MFs. RoT, instead, is used as part of the learning paradigm. RoT is not selected as a learning objective because: (i) it does not have a fixed answer and, more crucially, multiple RoTs may exist in one situationWe refer readers to Section 3.1 of ziems2022moral for further details., and (ii) it has a rather dynamic relationship with MFs.
@@
Experimental Settings. Following the steps of our pragmatic inference frameworks (detailed in Section 4), we design several questions and prompt the deepseek model to answer them. The step-wise answers are then used as the training dataThe data used in our experiments is available in the anonymized GitHub repository: https://anonymous.4open.science/r/moralpragmatics-F891. for fine-tuning the LLMs towards moral reasoning objectives. We chose Llama3-3.2-1B, Llama3-3.2-3B grattafiori2024llama3herdmodels and Mistral-7B (Mistral-7B-v0.3 jiang2023mistral7b) as our backbone models, and leverage the supervised fine-tuning objective to train those models by following previous studies ziems2022moral; liu2025diagnosing. We intentionally select smaller models, as they are generally considered less capable than their larger counterparts. We aim to demonstrate the effectiveness of our pragmatic inference frameworks through the strong performance achieved by these smaller models.
@@
To reiterate, we have three pragmatic inference frameworks, corresponding to three tasks. For the moral judgment task (Agree, Neutral, or Disagree with the reply), we use prediction accuracy as the evaluation metric. For the moral foundations classification (MFC) task, we compute prediction accuracy according to the number of MFs. Specifically, accuracies are reported separately for cases where a single RoT corresponds to one MF and for cases referring to two or more MFs. In other words, we have taken into consideration the one-to-many relationship between RoT/situation and MFs. We denote the LLM $f$ is parameterized by $\theta$, and the definitions (see Figure 5) of MFs are represented with $d_{m}$. Let $(x_{S},x_{R},y_{M},y_{J})$ denote a sample from the MIC benchmark, where $x_{S},x_{R}$ represents the moral situation and RoT, respectively. $y_{M}$ is its associated MF label. Each $y_{M}$ may contain between one and six MFs. $y_{J}$ is the moral judgment. We use $|y_{M}|=i$ $(1\leq i\leq 6,x\in\mathbb{Z})$ to indicate that $x$ is associated with $i$ MFs. We compute accuracy for each value of $i$.
@@
Figure 1: Prompt Questions For the Task of Moral Foundations Classification. All content within { } is provided by the benchmark.
@@
For baseline approaches, we follow prior work forbes2020social; ziems2022moral and consider two settings: (1) The base setting. For the moral judgment task, we compute $y_{J}=f_{\theta}(x_{S})$; and for MFC, we compute $y_{M}=f_{\theta}(x_{R})$. (2) The base+ setting. For the moral judgment task, we compute $y_{J}=f_{\theta}(x_{S},y_{M})$; and for MFC task, we compute $y_{M}=f_{\theta}(x_{R},d_{m})$. We additionally include the definitions of the six MFs ($d_{m}$) in the prompts. When the objective is moral judgment, we also explicitly provide the ground-truth MFs in the prompts. In other words, we intentionally provide the models with all available information in order to build strong baselines, which are then used to compare with the results of our framework. The setting of our proposed method is in line with the base+ setting but we incorporate an additional inference process in the input, e.g., $y_{J}=f_{\theta}(x_{S},y_{M},\{\text{inference}\})$.
@@
To examine how moral reasoning performance scales with the size of the training corpus, we consider three data settings. The benchmark reports agreement levels, and 26% of the samples receive full agreement between all annotators. We use subsets of 5K, 10K, and all of the 23500 samples of the full-agreement data. All performance results are reported based on evaluation using the development set. More details about hyperparameters and experimental setup are in Appendix A.1.
@@ 4 Pragmatic Inference
In this section, we introduce our pragmatic inference methods for different moral reasoning objectives. Please note, pragmatic inferences often have more than one pathway to achieve a conclusion. Our approach represents one such pathway. Specifically, we build three inferential methods for moral foundation classification, moral judgment based on ground-truth moral foundations, and the joint task of moral foundations classification and moral judgment, respectively.
@@ 4.1 Motivation
Proposing a pragmatic inference method for moral reasoning is challenging because there is no linguistic theory that clearly defines what constitutes an appropriate pragmatic inference for LLMs. Although several theories describe how humans make moral judgments, LLMs lack human-like linguistic competence bender-koller-2020-climbing; mahowald2024dissociating, making these theories unsuitable for direct application. On the other hand, LLMs are well known to capture distributional knowledge (i.e., distributional semantics) through statistical associations, and prior works have shown that they cannot leverage this distributional knowledge for satisfactory moral reasoning performance liu-etal-2025-diagnosing; lenci2023distributional. However, the distributional knowledge is much beyond distributional semantics, this is because pre-training corpora also contains static and context-limited texts graesser1994constructing; sap-etal-2022-neural, forcing LLMs not only capture distributional semantics but also approximate pragmatic cues through shallow statistical correlations.
For instance, the statement “calling your uncle by his given name is rude” is acceptable in many East Asian cultures but may not be perceived the same way elsewhere. Because pre-training texts provide no explicit contextual cues about the cultural setting in which such statements are embedded, LLMs may learn this norm only by associating negative descriptors “is rude” with the action “calling your uncle by his name” even though they lack any actual cultural understanding pawar2025survey. This implies that LLMs can leverage these superficial associations to predict certain pragmatic cues. However, the extent to which they can do so remains unknown, and we cannot rely on a purely semantics-driven approach to simplify the pragmatic inference. We hypothesize that MFT-driven pragmatic inference can help LLMs address the challenges outlined above, based on the following two perspectives.
@@
From the linguistic standpoint, our design of pragmatic inference methods is guided by three principles: (1) MFs are common ground which underlie all interactions. This has been supported by the theory and experiments of MFs graham2013moral; abdulhai2024moral. (2) MFs are abstract values extracted from discursive situations and hence need to be reified as context-specific interpretations graham_mapping_2011. It is these context-specific interpretations of MFs which provide effective semantic cues for the models to reason morally. (3) Moral judgment is the judgment of good or bad by its nature, more specifically, the good/bad evaluation placed on the outcome of an action nozick1968moral; schaich2006consequences; emelin2021moral. Therefore, it is important to build the connections between contexts, discursive interpretations of MFs, and the outcome (i.e., behavioral consequence).
From a statistical standpoint, liu2025diagnosing demonstrate that the pragmatic gap between moral situations and moral reasoning objectives introduces greater degradation in language modeling performance than purely semantics-driven tasks. This is because LLMs tend to develop spurious associations, between moral situations and reasoning objectives, that undermine distributional semantics. To address this, MFT provides a principled signal: all moral situations can be interpreted through its lens. The key, then, is to leverage our pragmatic inference methods to supply linguistic cues that appropriately map moral situations to the LLMs’ hidden state space that is characterized by MFs, thereby alleviating the degradation of distributional semantics.
Figure 2: Prompting Questions for the Task of Moral Judgment Based on Ground-truth Moral Foundations. All content within {} is provided by the benchmark.